OpenStudy (anonymous):

rewrite without rational exponents 27^4/3

ganeshie8 (ganeshie8):

Hint : \(\large 27=3^3\)

ganeshie8 (ganeshie8):

\[\large 27^{4/3}=(3^3)^{4/3}=3^{3*4/3}=3^4=?\]